# Noise-induced resonant acceleration of a charge in an intermittent magnetic field: an exact solution for ergodic and non-ergodic fluctuations

### v1

@@ V.1.2 Evaluation of $\langle r^{2}(t)\rangle$: the case of non-ergodic and ergodic power law fluctuations
$\langle r^{2}(s)\rangle=\frac{N(s)}{D(s)}$ (S20)
@@
$\begin{split}&D(s)=s(1+4s^{2})(9+4s^{2})(1+\sqrt{s\tau})(2+\sqrt{2}\sqrt{(-i+2s)\tau})(2+\sqrt{2}\sqrt{(i+2s)\tau})(2+\sqrt{2}\sqrt{(-3i+2s)\tau})(2+\sqrt{2}\sqrt{(3i+2s)\tau})\\ &\times\bigg(2\sqrt{9+4s^{2}}\sqrt{2+8s^{2}}\tau^{2}(-8+9\sqrt{s\tau})+4\tau\bigg(17\sqrt{2+8s^{2}}\sqrt{s\tau}+17\sqrt{18+8s^{2}}\sqrt{s\tau}-4\sqrt{9+4s^{2}}\sqrt{(-i+2s)\tau}\\ &+13\sqrt{9+4s^{2}}\sqrt{s\tau}\sqrt{(-i+2s)\tau}-4\sqrt{9+4s^{2}}\sqrt{(i+2s)\tau}+13\sqrt{9+4s^{2}}\sqrt{s\tau}\sqrt{(i+2s)\tau}-4\sqrt{1+4s^{2}}\sqrt{(-3i+2s)\tau}\\ &+13\sqrt{1+4s^{2}}\sqrt{s\tau}\sqrt{(-3i+2s)\tau}-4\sqrt{1+4s^{2}}\sqrt{(3i+2s)\tau}+13\sqrt{1+4s^{2}}\sqrt{s\tau}\sqrt{(3i+2s)\tau}\bigg)+4\sqrt{2}\sqrt{s\tau}\bigg(\sqrt{(-i+2s)\tau}\sqrt{(-3i+2s)\tau}\\ &+16\sqrt{(i+2s)\tau}\sqrt{(-3i+2s)\tau}+16\sqrt{(-i+2s)\tau}\sqrt{(3i+2s)\tau}+\sqrt{(i+2s)\tau}\sqrt{(3i+2s)\tau}\bigg)+s\tau\bigg(9\sqrt{9+4s^{2}}\sqrt{2+8s^{2}}\tau^{2}\\ &+18\tau\bigg(\sqrt{2+8s^{2}}+\sqrt{18+8s^{2}}+\sqrt{9+4s^{2}}\sqrt{(-i+2s)\tau}+\sqrt{9+4s^{2}}\sqrt{(i+2s)\tau}+\sqrt{1+4s^{2}}\sqrt{(-3i+2s)\tau}+\sqrt{1+4s^{2}}\sqrt{(3i+2s)\tau}\bigg)\\ -&2\bigg(16\sqrt{(-i+2s)\tau}+16\sqrt{(i+2s)\tau}+16\sqrt{(-3i+2s)\tau}+7\sqrt{2}\sqrt{(-i+2s)\tau}\sqrt{(-3i+2s)\tau}-8\sqrt{2}\sqrt{(i+2s)\tau}\sqrt{(-3i+2s)\tau}\\ &+16\sqrt{(3i+2s)\tau}-8\sqrt{2}\sqrt{(-i+2s)\tau}\sqrt{(3i+2s)\tau}+7\sqrt{2}\sqrt{(i+2s)\tau}\sqrt{(3i+2s)\tau}\bigg)\bigg)\bigg)\end{split}$ (S21)
we can plot the denominator $D(s)$ which shows a pole at $s\approx 0.0125$ (see Fig. 6) which coincides with the slope of the black line in Fig. 3 in main text.
@@
$\hat{\psi}(s)=\frac{\alpha}{\tau}e^{s\tau}E_{\alpha+1}[s]$ (S22)
@@
$\langle r^{2}(s)\rangle=\frac{N(s)}{D(s)}$ (S23)
@@
$\begin{split}&D(s)=3s(1+4s^{2})(9+4s^{2})\times\\ &\biggl(\text{E}_{5/2}[s]^{2}\biggl(-400e^{2i+2s}+36e^{i+4s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\Bigl(\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]+16e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\\ &\qquad-9e^{4s}\text{E}_{5/2}\Bigl[-\frac{i}{2}+s\Bigr]\biggl(-4e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]+\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]\Bigl(-64+81e^{2i+2s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\biggr)\biggr)\\ &\qquad+4\biggl(16e^{2i}-4e^{i+2s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\Bigl(\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]+16e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\\ &\qquad+e^{2s}\text{E}_{5/2}\Bigl[-\frac{i}{2}+s\Bigr]\biggl(-4e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]+\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]\Bigl(-64+225e^{2i+2s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\biggr)\biggr)\\ &\qquad\times\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]-32e^{\frac{i}{2}+2s}\text{E}_{5/2}[s]\biggl(-4\Bigl(\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]+e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\\ &\qquad+e^{2i}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\Bigl(-4+9e^{2s}\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)+\text{E}_{5/2}\Bigl[-\frac{i}{2}+s\Bigr]\biggl(9e^{2s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\\ &\qquad\times\Bigl(\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]+e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)+e^{i}\Bigl(-4+9e^{2s}\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\biggr)\biggr)\biggr)\end{split}$ (S24)
@@
Figure S2: Denominators for the mean square radius for $\omega_{0}=1$ and $\omega_{1}=0.5$ for the case of Mittag-Leffler function (left, with $\alpha=1/2$ and $\tau=10$) and power-law distribution with $\alpha=3/2$ and $\tau=1$ (right) used in main text. Both denominators show a single positive pole which determines the exponential rate of the mean radius square, these are the slopes of the black lines in Fig. 3 in main text.
V.2 DENSITY APPROACH
V.2.1 Evaluation of $\langle z(t)\rangle$
@@
$\frac{\partial^{2}U}{\partial t^{2}}+i\omega_{0}\frac{\partial U}{\partial t}+iW_{p}-\frac{1}{2}i\gamma W=0$ (S25)
$\frac{\partial W}{\partial t}=W_{p}-\gamma W$ (S26)
@@
$-i\omega_{1}^{2}\frac{\partial U}{\partial t}-\left(\gamma+i\omega_{0}\right)W_{p}$ (S27)
where we used that $\langle\dot{\omega}_{1}z\rangle=-\gamma W$, $\langle\dot{\omega}_{1}\dot{z}\rangle=-\gamma W_{p}$, and we used Eq. (5) for $\ddot{z}$. Taking the derivative of Eq. (S25) jointly with Eq. (S25) itself, we are able to find $W$ and $W_{p}$ as a function of $U$ and its derivatives. Taking a further derivative, we end up with
@@
$\gamma\left(i\gamma\omega_{0}-\omega_{0}^{2}+\omega_{1}^{2}\right)U^{\prime}(t)+\frac{1}{4}\gamma^{2}\omega_{1}^{2}U(t)=0$ (S28)
@@
$U(0)=0,\,\,U^{\prime}(0)=iv_{0},\,\,U^{\prime\prime}(0)=\omega_{0}v_{0},\,\,U^{\prime\prime\prime}(0)=-i\left(\omega_{0}^{2}+\omega_{1}^{2}\right)v_{0}$ (S29)
as can be inferred from Eqs. (S25)-(S27). The characteristic equation for Eq. (S28) is
$\lambda^{4}+2\left(\gamma+i\omega_{0}\right)\lambda^{3}+\left(\gamma^{2}+3i\gamma\omega_{0}-\omega_{0}^{2}+\omega_{1}^{2}\right)\lambda^{2}+\gamma\left(i\gamma\omega_{0}-\omega_{0}^{2}+\omega_{1}^{2}\right)\lambda+\frac{1}{4}\gamma^{2}\omega_{1}^{2}=0,$ (S30)
(which coincides with Eq. (S16)) and the solution can be found as
$U(t)=\sum_{i=1}^{4}c_{i}\exp[\lambda_{i}t]$ (S31)
with $c_{i}$ constants determined by conditions (S29), and $\lambda_{i}$ solutions of Eq. (S30).
V.2.2 Evaluation of $\langle r^{2}(t)\rangle$
@@
$r^{(6)}_{2}+3\gamma r^{(5)}_{2}+\left[3\gamma^{2}+2\left(\omega_{0}^{2}+\omega_{1}^{2}\right)\right]r^{(4)}_{2}+\gamma\left[\gamma^{2}+4\left(\omega_{0}^{2}+\omega_{1}^{2}\right)\right]r^{(3)}_{2}$+ (S32)
@@
The characteristic equation for Eq. (V.2.2) is
$\lambda^{6}+3\gamma\lambda^{5}+\left[3\gamma^{2}+2\left(\omega_{0}^{2}+\omega_{1}^{2}\right)\right]\lambda^{4}_{2}+\gamma\left[\gamma^{2}+4\left(\omega_{0}^{2}+\omega_{1}^{2}\right)\right]\lambda^{3}+$ (S33)
@@
Both Eqs. (S28) and (V.2.2) can be solved analytically, being differential equations with constant coefficients. Being the characteristic equations of the fourth and sixth degree, respectively, the expressions for the constants $\lambda_{i}$ are not easy to handle.
Let us focus on the solution of Eq. (V.2.2), describing the diffusion of the particle. Using Descartes’ rule of signs, we deduce that there is a positive root, which implies that, in general, the particle diffuses and $r_{2}(t)\to\infty$ for $t\to\infty$. The full solution of Eq. (V.2.2) gives an unmanageable analytical expression. We may find an analytical approximated solution considering the case $\omega_{0}\ll\omega_{1}$. Due to the evident symmetry of the polynomial with respect to the exchange of $\omega_{0}$ and $\omega_{1}$, the result will also apply to $\omega_{1}\ll\omega_{0}$. Setting $\varepsilon\equiv\omega_{0}/\omega_{1}$ we have
$\lambda^{6}+3\gamma\lambda^{5}+\left[3\gamma^{2}+2\omega_{1}^{2}\left(1+\varepsilon^{2}\right)\right]\lambda^{4}+\gamma\left[\gamma^{2}+4\omega_{1}^{2}\left(1+\varepsilon^{2}\right)\right]\lambda^{3}+$ (S34)
@@
$\lambda_{1}=\varepsilon^{2}\frac{\gamma\omega_{1}^{2}}{\gamma^{2}+\omega_{1}^{2}},$ (S35)
$\lambda_{2}=-\gamma-\varepsilon^{2}\frac{\gamma\omega_{1}^{2}}{\gamma^{2}+\omega_{1}^{2}},$ (S36)
$\lambda_{3}=-\gamma+i\omega_{1}+\varepsilon^{2}\left[\frac{\omega_{1}^{2}}{\gamma}+\frac{1}{2}\frac{i\gamma\omega_{1}}{\gamma-i\omega_{1}}\right],$ (S37)
$\lambda_{4}=-\gamma-i\omega_{1}+\varepsilon^{2}\left[\frac{\omega_{1}^{2}}{\gamma}-\frac{1}{2}\frac{i\gamma\omega_{1}}{\gamma+i\omega_{1}}\right],$ (S38)
$\lambda_{5}=i\omega_{1}+\varepsilon^{2}\left[\frac{1}{2}\frac{i\gamma\omega_{1}}{\gamma+i\omega_{1}}-\frac{\omega_{1}^{2}}{\gamma}\right]$ (S39)
$\lambda_{6}=-i\omega_{1}+\varepsilon^{2}\left[-\frac{1}{2}\frac{i\gamma\omega_{1}}{\gamma-i\omega_{1}}-\frac{\omega_{1}^{2}}{\gamma}\right]$ (S40)

### v3

@@ VI.0.2 Evaluation of $\langle r^{2}(t)\rangle$: the case of non-ergodic and ergodic power law fluctuations
$\langle r^{2}(s)\rangle=\frac{N(s)}{D(s)}$ (A20)
@@
$\begin{split}&D(s)=s(1+4s^{2})(9+4s^{2})(1+\sqrt{s\tau})(2+\sqrt{2}\sqrt{(-i+2s)\tau})(2+\sqrt{2}\sqrt{(i+2s)\tau})(2+\sqrt{2}\sqrt{(-3i+2s)\tau})(2+\sqrt{2}\sqrt{(3i+2s)\tau})\\ &\times\bigg(2\sqrt{9+4s^{2}}\sqrt{2+8s^{2}}\tau^{2}(-8+9\sqrt{s\tau})+4\tau\bigg(17\sqrt{2+8s^{2}}\sqrt{s\tau}+17\sqrt{18+8s^{2}}\sqrt{s\tau}-4\sqrt{9+4s^{2}}\sqrt{(-i+2s)\tau}\\ &+13\sqrt{9+4s^{2}}\sqrt{s\tau}\sqrt{(-i+2s)\tau}-4\sqrt{9+4s^{2}}\sqrt{(i+2s)\tau}+13\sqrt{9+4s^{2}}\sqrt{s\tau}\sqrt{(i+2s)\tau}-4\sqrt{1+4s^{2}}\sqrt{(-3i+2s)\tau}\\ &+13\sqrt{1+4s^{2}}\sqrt{s\tau}\sqrt{(-3i+2s)\tau}-4\sqrt{1+4s^{2}}\sqrt{(3i+2s)\tau}+13\sqrt{1+4s^{2}}\sqrt{s\tau}\sqrt{(3i+2s)\tau}\bigg)+4\sqrt{2}\sqrt{s\tau}\bigg(\sqrt{(-i+2s)\tau}\sqrt{(-3i+2s)\tau}\\ &+16\sqrt{(i+2s)\tau}\sqrt{(-3i+2s)\tau}+16\sqrt{(-i+2s)\tau}\sqrt{(3i+2s)\tau}+\sqrt{(i+2s)\tau}\sqrt{(3i+2s)\tau}\bigg)+s\tau\bigg(9\sqrt{9+4s^{2}}\sqrt{2+8s^{2}}\tau^{2}\\ &+18\tau\bigg(\sqrt{2+8s^{2}}+\sqrt{18+8s^{2}}+\sqrt{9+4s^{2}}\sqrt{(-i+2s)\tau}+\sqrt{9+4s^{2}}\sqrt{(i+2s)\tau}+\sqrt{1+4s^{2}}\sqrt{(-3i+2s)\tau}+\sqrt{1+4s^{2}}\sqrt{(3i+2s)\tau}\bigg)\\ -&2\bigg(16\sqrt{(-i+2s)\tau}+16\sqrt{(i+2s)\tau}+16\sqrt{(-3i+2s)\tau}+7\sqrt{2}\sqrt{(-i+2s)\tau}\sqrt{(-3i+2s)\tau}-8\sqrt{2}\sqrt{(i+2s)\tau}\sqrt{(-3i+2s)\tau}\\ &+16\sqrt{(3i+2s)\tau}-8\sqrt{2}\sqrt{(-i+2s)\tau}\sqrt{(3i+2s)\tau}+7\sqrt{2}\sqrt{(i+2s)\tau}\sqrt{(3i+2s)\tau}\bigg)\bigg)\bigg)\end{split}$ (A21)
we can plot the denominator $D(s)$ which shows a pole at $s\approx 0.0125$ (see Fig. A2) which coincides with the slope of the black line in Fig. 3 in main text.
@@
$\hat{\psi}(s)=\frac{\alpha}{\tau}e^{s\tau}E_{\alpha+1}[s]$ (A22)
@@
$\langle r^{2}(s)\rangle=\frac{N(s)}{D(s)}$ (A23)
@@
$\begin{split}&D(s)=3s(1+4s^{2})(9+4s^{2})\times\\ &\biggl(\text{E}_{5/2}[s]^{2}\biggl(-400e^{2i+2s}+36e^{i+4s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\Bigl(\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]+16e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\\ &\qquad-9e^{4s}\text{E}_{5/2}\Bigl[-\frac{i}{2}+s\Bigr]\biggl(-4e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]+\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]\Bigl(-64+81e^{2i+2s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\biggr)\biggr)\\ &\qquad+4\biggl(16e^{2i}-4e^{i+2s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\Bigl(\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]+16e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\\ &\qquad+e^{2s}\text{E}_{5/2}\Bigl[-\frac{i}{2}+s\Bigr]\biggl(-4e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]+\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]\Bigl(-64+225e^{2i+2s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\biggr)\biggr)\\ &\qquad\times\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]-32e^{\frac{i}{2}+2s}\text{E}_{5/2}[s]\biggl(-4\Bigl(\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]+e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\\ &\qquad+e^{2i}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\Bigl(-4+9e^{2s}\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)+\text{E}_{5/2}\Bigl[-\frac{i}{2}+s\Bigr]\biggl(9e^{2s}\text{E}_{5/2}\Bigl[\frac{i}{2}+s\Bigr]\\ &\qquad\times\Bigl(\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]+e^{3i}\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)+e^{i}\Bigl(-4+9e^{2s}\text{E}_{5/2}\Bigl[-\frac{3i}{2}+s\Bigr]\text{E}_{5/2}\Bigl[\frac{3i}{2}+s\Bigr]\Bigr)\biggr)\biggr)\biggr)\end{split}$ (A24)
@@
Figure A2: Denominators for the mean square radius for $\omega_{0}=1$ and $\omega_{1}=0.5$ for the case of Mittag-Leffler function (left, with $\alpha=1/2$ and $\tau=10$) and power-law distribution with $\alpha=3/2$ and $\tau=1$ (right) used in main text. Both denominators show a single positive pole which determines the exponential rate of the mean radius square, these are the slopes of the black lines in Fig. 3 in main text.
VII S3: DENSITY APPROACH
VII.0.1 Evaluation of $\langle z(t)\rangle$
@@
$\frac{\partial^{2}U}{\partial t^{2}}+i\omega_{0}\frac{\partial U}{\partial t}+iW_{p}-\frac{1}{2}i\gamma W=0$ (A25)
$\frac{\partial W}{\partial t}=W_{p}-\gamma W$ (A26)
@@
$-i\omega_{1}^{2}\frac{\partial U}{\partial t}-\left(\gamma+i\omega_{0}\right)W_{p}$ (A27)
where we used that $\langle\dot{\omega}_{1}z\rangle=-\gamma W$, $\langle\dot{\omega}_{1}\dot{z}\rangle=-\gamma W_{p}$, and we used Eq. (5) for $\ddot{z}$. Taking the derivative of Eq. (A25) jointly with Eq. (A25) itself, we are able to find $W$ and $W_{p}$ as a function of $U$ and its derivatives. Taking a further derivative, we end up with
@@
$\gamma\left(i\gamma\omega_{0}-\omega_{0}^{2}+\omega_{1}^{2}\right)U^{\prime}(t)+\frac{1}{4}\gamma^{2}\omega_{1}^{2}U(t)=0$ (A28)
@@
$U(0)=0,\,\,U^{\prime}(0)=iv_{0},\,\,U^{\prime\prime}(0)=\omega_{0}v_{0},\,\,U^{\prime\prime\prime}(0)=-i\left(\omega_{0}^{2}+\omega_{1}^{2}\right)v_{0}$ (A29)
as can be inferred from Eqs. (A25)-(A27). The characteristic equation for Eq. (A28) is
$\lambda^{4}+2\left(\gamma+i\omega_{0}\right)\lambda^{3}+\left(\gamma^{2}+3i\gamma\omega_{0}-\omega_{0}^{2}+\omega_{1}^{2}\right)\lambda^{2}+\gamma\left(i\gamma\omega_{0}-\omega_{0}^{2}+\omega_{1}^{2}\right)\lambda+\frac{1}{4}\gamma^{2}\omega_{1}^{2}=0,$ (A30)
(which coincides with Eq. (A16)) and the solution can be found as
$U(t)=\sum_{i=1}^{4}c_{i}\exp[\lambda_{i}t]$ (A31)
with $c_{i}$ constants determined by conditions (A29), and $\lambda_{i}$ solutions of Eq. (A30).
VII.0.2 Evaluation of $\langle r^{2}(t)\rangle$
@@
$r^{(6)}_{2}+3\gamma r^{(5)}_{2}+\left[3\gamma^{2}+2\left(\omega_{0}^{2}+\omega_{1}^{2}\right)\right]r^{(4)}_{2}+\gamma\left[\gamma^{2}+4\left(\omega_{0}^{2}+\omega_{1}^{2}\right)\right]r^{(3)}_{2}$+ (A32)
@@
The characteristic equation for Eq. (VII.0.2) is
$\lambda^{6}+3\gamma\lambda^{5}+\left[3\gamma^{2}+2\left(\omega_{0}^{2}+\omega_{1}^{2}\right)\right]\lambda^{4}_{2}+\gamma\left[\gamma^{2}+4\left(\omega_{0}^{2}+\omega_{1}^{2}\right)\right]\lambda^{3}+$ (A33)
@@
Both Eqs. (A28) and (VII.0.2) can be solved analytically, being differential equations with constant coefficients. Being the characteristic equations of the fourth and sixth degree, respectively, the expressions for the constants $\lambda_{i}$ are not easy to handle.
Let us focus on the solution of Eq. (VII.0.2), describing the diffusion of the particle. Using Descartes’ rule of signs, we deduce that there is a positive root, which implies that, in general, the particle diffuses and $r_{2}(t)\to\infty$ for $t\to\infty$. The full solution of Eq. (VII.0.2) gives an unmanageable analytical expression. We may find an analytical approximated solution considering the case $\omega_{0}\ll\omega_{1}$. Due to the evident symmetry of the polynomial with respect to the exchange of $\omega_{0}$ and $\omega_{1}$, the result will also apply to $\omega_{1}\ll\omega_{0}$. Setting $\varepsilon\equiv\omega_{0}/\omega_{1}$ we have
$\lambda^{6}+3\gamma\lambda^{5}+\left[3\gamma^{2}+2\omega_{1}^{2}\left(1+\varepsilon^{2}\right)\right]\lambda^{4}+\gamma\left[\gamma^{2}+4\omega_{1}^{2}\left(1+\varepsilon^{2}\right)\right]\lambda^{3}+$ (A34)
@@
$\lambda_{1}=\varepsilon^{2}\frac{\gamma\omega_{1}^{2}}{\gamma^{2}+\omega_{1}^{2}},$ (A35)
$\lambda_{2}=-\gamma-\varepsilon^{2}\frac{\gamma\omega_{1}^{2}}{\gamma^{2}+\omega_{1}^{2}},$ (A36)
$\lambda_{3}=-\gamma+i\omega_{1}+\varepsilon^{2}\left[\frac{\omega_{1}^{2}}{\gamma}+\frac{1}{2}\frac{i\gamma\omega_{1}}{\gamma-i\omega_{1}}\right],$ (A37)
$\lambda_{4}=-\gamma-i\omega_{1}+\varepsilon^{2}\left[\frac{\omega_{1}^{2}}{\gamma}-\frac{1}{2}\frac{i\gamma\omega_{1}}{\gamma+i\omega_{1}}\right],$ (A38)
$\lambda_{5}=i\omega_{1}+\varepsilon^{2}\left[\frac{1}{2}\frac{i\gamma\omega_{1}}{\gamma+i\omega_{1}}-\frac{\omega_{1}^{2}}{\gamma}\right]$ (A39)
$\lambda_{6}=-i\omega_{1}+\varepsilon^{2}\left[-\frac{1}{2}\frac{i\gamma\omega_{1}}{\gamma-i\omega_{1}}-\frac{\omega_{1}^{2}}{\gamma}\right]$ (A40)